The midpoints of the sides of a triangle are at (0,-3) , (-1/2,7/2) and (-7/2,1/2) Find the coordinates of the vertices.
Answer by mananth(16949)

The midpoints of the sides of a triangle are at (0,-3) , (-1/2,7/2) and (-7/2,1/2) Find the coordinates of the vertices.
.
By mid point formula (x1+x2)/2 = 0 ;x1+x2=0............(1)
(x2+x3)/2 = -1/2 ; x2+x3=-1............(2)
(x1+x3)/2 = -7/2 ; x1+x3= -7.............(3)
substitute x2 in (1)
x1+(-1-x3) = 0
x1-x3=1.....(3)
solve with
x1+x3=-7.......(4)
add (3) & (4)
2x1=-6
x1=-3
x1+x3=-7
-3+x3=-7
x3= -4
x1+x2=0
x2=3
(x1=-3,x2=3,x3=-4)
Similarly solve for y1,y2,y3